A student appear a test of 5 subjects.how many different ways he can be failed
Answer by mathmate(429)


Question:
A student appear a test of 5 subjects.how many different ways he can be failed

Solution:
Order is not important here.
For the first subject, there are two outcomes, pass or fail.
The same number of outcomes for each of the following subjects.
Consider this a 5 step experiment, each with two outcomes.
Using the multiplication rule, the total number of possible results is
2*2*2*2*2=2^5=32.
To generalize, with n subjects, there are 2^n results.